Solve by factoring:
x^2-9x=-20
Thanks in advance!
C

Answer by Caallen2(9)

This one is solved the same as the previous one that I submitted to you.
x^2 - 9x = -20
simplify
x^2 - 9x + 20 = 0
then form two simpler equations
(x - 4)(x - 5)= 0
x = 4 x = 5